solve by the substitution method.
6m+n= 28
m-7n=62 what is the solution I came up with ( 4,-8) am I close

Answer by Robb(9)

You got one right!!! Good job.
Here is the problem worked for you:
I will solve 6M + N = 28 for "M"
6M + N = 28
6M = 28 - N
M = (28 - N)/6
Now that we have that answer for M, substitute it into the equation M - 7N = 62 to give us this equation:
(28 - N)/6 - 7N = 62
First thing we want to do now is get rid of our fraction!
We do that by multiplying each of the 3 terms by the denominator 6. That will give us this equation:
28 - N - 42N = 372
Now we combine like terms:
28 - 43N = 372
Next, we subtract 28 from each side of the equation:
****don't forget that leaves us with a NEGATIVE 43N****
-43N = 344
Finally, we divide both sides by a -43 to get:
N=-8
Now that we have N's value, we plug that back into either original equation you want, I'll use 6M + N = 28
Plugging N's value in to that equation, we get:
6M + (-8) = 28
Subtract (-8) from both sides:
***subtacting a negative number turns it to a positive***
Now we have:
6M = 36
Divide each side by 6 to get:
M = 6
So your answer is M=6 and N=-8
You came close, probably just a small arithmatic mistake somewhere...
Hope this helps!